OpenStudy (anonymous):

Germain ate 1/4 of a bag of popcorn that was 1/2 full. How much of the whole bag of popcorn did she eat? A. 1/4bag B.1/2 bag C. 18/ bag D. 3/4 bag

OpenStudy (ilovecake):

I just did this test! Today!

OpenStudy (jdoe0001):

he ate 1/4 of a bag that was 1/2 already so he really ate 1/2 of 1/4 or \(\bf \cfrac{1}{4}\times \cfrac{1}{2}\)

OpenStudy (ilovewolf):

1/4 * 1/2 = ? = ( 1* 1) / ( 4*2) = 1/8 So the answer would be C c:

OpenStudy (ilovecake):

I took the test and it was 18

OpenStudy (anonymous):

Really @Ilovecake do you know the rest? I have 3 more

OpenStudy (ilovewolf):

18? or 1/8?

OpenStudy (ilovecake):

Okay

OpenStudy (ilovecake):

18

OpenStudy (ilovewolf):

oh okay c:

OpenStudy (ilovecake):

I know the rest